# Strong Coupling Constant Determination from the new CTEQ-TEA Global QCD Analysis

### v1

@@ III.1 The PDF parametrization uncertainty is small…
where the 68% CL is computed according to $\Delta\chi^{2}=37$, the increase in the global $\chi^{2}$ consistent with the CT18 tolerance definition. This $\alpha_{s}(M_{Z})$ value is in a good agreement with the current world average. At the same time, we observe that the preferred $\alpha_{s}(M_{Z})$ values vary across the processes. For example, the DIS experiments prefer the lowest value ($\alpha_{s}(M_{Z})=0.1152$). Drell–Yan pair and top-quark pair production prefer higher values, 0.1198 and 0.1194, respectively. Inclusive jet production prefers an intermediate value ($\alpha_{s}(M_{Z})=0.1187$). The spread of preferences in Fig. 4 (left) reflects the differing pulls on $\alpha_{s}(M_{Z})$ from various experiments and highlights tensions in the data.
When the ksys=2 prescription is applied to the jet data in Fig. 4 (right), the global minimum of total $\chi^{2}_{R}$ prefers a much lower $\alpha_{s}(M_{Z})$ value,
@@
Refitting the PDFs results in an even larger downward shift of the preferred $\alpha_{s}(M_{Z})$ with ksys=2 then when the PDFs were fixed in the previous exercise.
In this scenario, the ordering of preferred values is markedly different: jet production measurements favor by far the smallest value ($\alpha_{s}(M_{Z})=0.1099$), while $t\bar{t}$ production still pulls toward the largest value ($\alpha_{s}(M_{Z})=0.1191$), and the DIS ($\alpha_{s}(M_{Z})=0.1146$) and Drell–Yan process ($\alpha_{s}(M_{Z})=0.1187$) lying in-between. Compared to the ksys=1 case, the pull from jet-production measurements to a significantly lower value of $\alpha_{s}(M_{Z})$ is accompanied by an equally significant reduction in $\chi^{2}$, which reflects strong sensitivity of these measurements to the treatment of correlated systematics.
@@
To further investigate the differences between the ksys=1 and ksys=2 scenarios in the case of inclusive-jet production measurements, Fig. 5 separately illustrates the $\Delta\chi^{2}$ profiles of individual jet experiments. As before, the curves have $\Delta\chi^{2}=0$ at the best-fit $\alpha_{s}(M_{Z})$ of the respective global fit. The ksys=1 case is shown in the Fig. 5 (left), where the DØ Run-2 and ATLAS measurements at 7, 8, and 13 TeV collision energies prefer relatively large $\alpha_{s}(M_{Z})$, while the CDF Run-2 and CMS measurements at 7, 8, and 13 TeV prefer smaller values. Their interplay produces a combined $\alpha_{s}(M_{Z})$-minimum from jet production that is
@@
and is consistent with the world average.
By contrast, when the ksys=2 treatment is applied to jet data sets in Fig. 5 (right), all of them prefer a much smaller $\alpha_{s}(M_{Z})$. The CMS 8 TeV and CDF Run-2 data especially pull toward smaller $\alpha_{s}(M_{Z})$ values, compared to the preference of DØ and ATLAS for a somewhat larger $\alpha_{s}(M_{Z})$. This produces a much lower central value
@@
that is well below the world average, with a larger uncertainty.
In summary, the comparison between the two different treatments for the correlated systematic uncertainties in inclusive jet production, ksys=1 and ksys=2, demonstrates that these measurements are particularly sensitive to the choice of treatment. As we emphasized, requiring all correlated errors to be normalized according to the ksys=2 (additive) choice may exaggerate the downward pull on $\alpha_{s}(M_{Z})$ and in fact introduce a statistical bias. An analogous downward shift has been noted in the NNPDF precision determination [24] of $\alpha_{s}(M_{Z})$. Its final $\alpha_{s}(M_{Z})$ is quoted for the additive prescription (the analog of our ksys=2). Given that the additive vs. multiplicative nature of the provided correlated errors is largely undisclosed, and other treatments may involve biases as well, this exercise shows that details of the systematic error models play the central yet understated role in the extraction of $\alpha_{s}(M_{Z})$.
IV Tolerances and data clustering safety
@@
Even when theory perfectly agrees with data, the $\chi^{2}_{E}/N_{\textrm{pt}}$ ratios of individual experiments fluctuate around the expectation of one with a known probability. What is actually seen in all global fits is that the spread of $\chi^{2}_{E}/N_{\textrm{pt}}$ is wider than this normal distribution arising from purely stochastic fluctuations. One can demonstrate this enlargement of discrepancies by considering histograms of effective Gaussian variables $S_{E}$ [25] constructed from $\chi^{2}_{E}$. While, in the absence of tensions, the distribution of $S_{E}$ is expected to be consistent with the standard normal distribution $\mathcal{N}(0,1)$, the empirically obtained histograms of $S_{E}$ for the CT, MSHT, and NNPDF global fits are wider than $\mathcal{N}(0,1)$. This behavior raises the question of finding consistent procedures to combine the $\alpha_{s}(M_{Z})$ values from individual experiments and their corresponding uncertainties.
We mentioned that the naive application of the $\Delta\chi^{2}=1$ criterion to the total $\chi^{2}$ in the CT25 fit would result in $\alpha_{s}(M_{Z})=0.1184\pm 0.0004$, which is smaller than the uncertainty of the world average quoted in the PDG [61]. In this and next section, we will review several strategies to obtain more realistic error estimates:
@@
    Introducing tolerance on the PDF errors in Sec. IV.1; exploration of data clustering safety of such prescriptions in Sec. IV.3;
@@
    Reclustering of data sets in Sec. IV.4;
@@
    Bayesian model averaging and data reweighting in Sec. V.
These are just some possible techniques, which nevertheless give a sense of what modifications to the $\Delta\chi^{2}=1$ prescription one might expect with the realistic data in the PDF fit.
IV.1 Global and Dynamical Tolerance Criteria
Tolerance-based methods aim to account for fit imperfections such as tensions by modifying the $\Delta\chi^{2}=1$ criterion assigning an increase in $\chi^{2}$ to the 68% CL uncertainty. The simplest global tolerance [68] uniformly associates the 68% CL uncertainty with an increase $\Delta\chi^{2}=T^{2}>1$ for any combination of PDF parameters. This prescription may be too crude, as e.g. tensions among experiments are likely to constrain different combinations of PDF parameters in unsimilar ways. The dynamic tolerance leverages constraints from individual experiments to obtain $T^{2}$ values independently for each orthogonal direction in the PDF parameter space in the Hessian formalism.
Global Tolerance. A common way to introduce either the global or dynamic tolerance is to consider quantiles of the probability density $P(x)$ supported on $\infty<x<\infty$. The quantile $\xi(P,v)$ at cumulative probability $v$ is defined by
@@ IV.1 Global and Dynamical Tolerance Criteria
To define the global tolerance (GT) at the 68% CL (corresponding to $v=0.68$) for $\alpha_{s}(M_{Z})$, one might consider $P(x)$ for the $\chi^{2}$ distribution of the whole data set. Then, one determines $T^{2}$, or equivalently constructs the 68% CL interval $\alpha_{s,\min}\leq\alpha_{s}(M_{Z})\leq\alpha_{s,\max}$, from the quantile $\xi(\chi^{2}(N_{\textrm{pt}}),0.68)$ of the $\chi^{2}$ distribution with $N_{\textrm{pt}}$ degrees of freedom.
The CT [39, 69] and MSHT [70] groups implement variations of this generic procedure in part to account for the more-than-normal excess of the best-fit $\chi^{2}_{0}$ above its expectation value $N_{\textrm{pt}}$. For the purpose of this discussion, we will approximate both the CT and MSHT procedures by the following simplified prescription based on the effective Gaussian variable $S$.
$S$ represents a linearized $Z$ score, i.e., it is a function of $\chi^{2}$ that is distributed approximately according to $\mathcal{N}(0,1)$ and is easier to work with than the $\chi^{2}$ distribution with $N_{\textrm{pt}}$ degrees of freedom. For each $\chi^{2}$, the corresponding $S(\chi^{2},N_{\textrm{pt}})$ can be computed using either a simpler Fisher’s approximation [71],
@@
Dynamical Tolerance. When the quantiles in Eq.(11) are computed for each $\chi^{2}_{E}$ profile of the individual data sets, and then a single uncertainty range is constructed from their combination, we refer to it as the “dynamical tolerance uncertainty” [70, 25]. Suppose the global tolerance (GT) applied just to experiment $E$ renders the uncertainty range $\alpha_{s,\min}^{(E)}\leq\alpha_{s}(M_{Z})\leq\alpha_{s,\max}^{(E)}$. The dynamic tolerance (DT) interval for the full data set is constructed separately from the upper and lower boundaries for each experiment as
The two-tier CT18 uncertainty. The global and dynamical tolerance can be combined into a more complex constraint, such as the two-tier tolerance prescription used by CT10, CT14, and CT18 PDFs. In this prescription, the global tolerance with $T^{2}\sim 37$ is combined with the “tier-2” dynamical tolerance at the 68% CL. The goal is of the current study is to bypass this definition, which otherwise would require, in particular, to estimate the global $T^{2}$ through the quantile argument or by another means. The dynamic tolerance, on the other hand, depends on how the data are partitioned into “experiments”. Different partitions modify the DT uncertainties, which leads to the concept of “data clustering safety” that we will soon discuss.
@@
IV.2 Properties of the dynamic tolerance
The dynamical tolerance (DT) proved to be convenient when accounting for constraints from individual experiments on PDF uncertainties. That said, at the detailed level, the DT behaves more elaborately than the GT. These pecularities come to the forefront when one strives to elevate its precision.
@@ IV.2 Properties of the dynamic tolerance
DT intervals tend to be asymmetric around the central value given by the position of the global minimum for all experiments. This can be explicitly seen from Eqs. (14) and (15), showing that the DT intervals are determined from changes in the probabilities of two most constraining experiments $E$ with respect to the global $\chi^{2}$ minimum. [^4] In the complete the CT or MSHT implementations of the DT formulas, such behavior is achieved by applying the quantile to the ratio $S(\chi^{2}_{E})/S(\chi^{2}_{E,0})$ or $\chi^{2}_{E}/\chi^{2}_{E,0}$, not to $S(\chi^{2}_{E})$ or $\chi^{2}_{E}$ itself. For example, Fig. 6 compares the GT and DT uncertainties in the CT25 NNLO fit for credibility levels indicated on the horizontal axis. In this case, the GT is found from the quantile of the baseline data set. The DT is constructed from the quantiles of 49 contributing data sets and has obvious asymmetry above and below the central $\alpha_{s}(M_{Z})$.
@@
Figure 7 shows the same default GT and DT uncertainties for the 49 data sets as in Fig. 6. In addition, we recomputed the DT 10000 times, every time randomly combining the 49 data sets into 24 clusters in the left subfigure and into 3 clusters in the right one. For every CL value, we thus quantify the uncertainty on the DT uncertainty, communicated in the figure by filled blue bands containing 6800 instances of the reclustered DT intervals, as well as the dotted blue bands indicating the most extreme instances among the 10000 reclustered DT intervals. From Fig. 7, it is clear that the average length and uncertainty of the DT interval depend on reclustering. Fewer clusters bring the DT uncertainty closer to the GT uncertainty, corresponding in this example to a single global data cluster. Conversely, finer clustering reduces the DT uncertainty interval.
@@
Figure 8 summarizes this behavior by showing the DT uncertainty at 68% CL and 90% CL for the CT25prel baseline data comprised of 52 published data sets. In addition to the GT uncertainty, corresponding to one cluster containing all data, we show the DT uncertainty for 52 published data sets, as well as for alternative clusters detailed in Sec. IV.4.
The figure makes it obvious that the DT uncertainty changes depending on the clustering procedure. When data are split into a large number of clusters (1000 in Fig. 8), the DT uncertainty can be essentially zero, as it becomes very likely that a few outlier data points somewhere in the data sample will lead to a very tight DT constraint on $\alpha_{s}(M_{Z})$.
Figure 8: Global and dynamical tolerance uncertainties on $\alpha_{s}(M_{Z})$ at the 68% and 90% CL for various clustering procedures applied to the CT25prel baseline. DT uncertainties are estimated according to the default prescription, without and with including the sum $R^{2}$ of best-fit nuisance parameters, and optionally by reclustering the data into 3, 52, 100, and 1000 clusters as explained in Sec. IV.4. For the reclustered data, the DT errors are averaged over 200 alternative configurations.
IV.3 Data clustering safety of uncertainty estimates
The previous subsection points out that the DT uncertainty depends on how the data are clustered. The fitting groups are constructing the DT using the full published data sets as independent clusters, but other reasonable clustering choices are obviously possible, often leading to different DT intervals.
Such sensitivity would raise concern that the DT uncertainty estimate on a fundamental parameter is unsafe at some level with respect to the user-chosen data clustering prescription. In contrast, the GT is trivially clustering-safe.
The key point here is that reclustering of the data sets modifies the degree of tensions present in the data. For instance, in a version of the CT25prel fit to 52 data sets shown in Fig. 8, the GT estimate is $\alpha_{s}(M_{Z})=0.1179\pm 0.0024$, i.e. the error is practically symmetric. The default DT error is $0.1179^{+0.0024}_{-0.0011}$, showing a pronounced asymmetry. Upon reorganizing the data sets into 3 clusters, without breaking the data sets, in some cases the data sets are grouped so that their opposite pulls on $\alpha_{s}(M_{Z})$, and in other cases these pulls are maximally exacerbated. When the tensions are largely cancelled, the 3-cluster DT error becomes nearly symmetric and larger than the GT one, e.g., $0.1179^{+0.0032}_{-0.0028}$. When the clustering magnifies the tensions, the 3-cluster DT errors vary more and can be highly asymmetric or not, for instance, $0.1179^{+0.0021}_{-0.0008}$ or $0.1179^{+0.0016}_{-0.0019}$.
The DT estimates thus in principle should be supplemented by additional tests of the degree of their clustering (un)safety. In light of what was said above, the DT errors for the $\alpha_{s}$ determination vary between zero and the GT errors. One could take an envelope of these DT errors as a conservative estimate, or an average DT error over many clustering configurations as an optimal estimate. In our specific case, the latter estimate is close to the default one, and hence the final uncertainty combination will concentrate on the default DT tolerance.
IV.4 Data reclustering and $K$-folding
In the remainder of this section, we will summarize the procedure employed to divide the published data sets into many clusters illustrated in Fig. 8. At the end, we will comment on the potential impact of cross validation procedures in the Monte-Carlo-based approaches on the $\alpha_{s}$ uncertainty estimate.
IV.4.1 Reclustering algorithm
Reclustering is based on the representation of the best-fit $\chi^{2}$ in Eq. (4) by the quadrature sums $D^{2}$ and $R^{2}$ of data residuals $r_{0,k}^{(E)}$ and nuisance parameters $\lambda^{(E)}_{0,\alpha}$. If we neglect correlations via the PDF parameters, as was done to justify the DT to start with, then $r_{0,k}^{(E)}]^{2}$ and $\lambda^{(E)}_{0,\alpha}]^{2}$ from all data sets can be separately and independently collected (“shuffled”) into alternative clusters of data points, and then the DT can be formed from such clusters.
To examine how the DT bands change when we partition the published data sets instead of combining them, we implement the following exercise. We work with the CT25prel baseline, create a grid of $N_{\alpha_{s}}=21$ values of $\alpha_{s}(M_{Z})$ between 0.113 and 0.129, and select a reference $\alpha_{s}(M_{Z})_{\textrm{ref}}$, which in this case we set equal to the best-fit $\alpha_{s}(M_{Z})=0.1179$. For each $\alpha_{s}$, we can organize all $\chi_{R}^{2}$ components in Eq. (4) into two vectors $\vec{v}_{r}$ and $\vec{v}_{\lambda}$ as
$\vec{v}_{r}$ $=$ $\left\{r^{(1)}_{0,1},\dots r^{(1)}_{0,N_{\textrm{pt}}^{(1)}},\dots r^{(N_{E})}_{0,1},\dots,r^{(N_{E})}_{0,{N_{\textrm{pt}}^{(N_{E}})}}\right\}\,$, (16)
$\vec{v}_{\lambda}$ $=$ $\left\{\lambda^{(1)}_{0,1},\dots,\lambda^{(1)}_{0,N_{\lambda}^{(1)}},\dots,\lambda^{(N_{E})}_{0,1},\dots,\lambda^{(N_{E})}_{0,N_{\lambda}^{(N_{E})}}\right\}\,$, (17)

### v2

@@ III.1 The PDF parametrization uncertainty is small…
where the 68% CL is computed according to $\Delta\chi^{2}=37$, the increase in the global $\chi^{2}$ consistent with the CT18 tolerance definition. This $\alpha_{s}$ value is in good agreement with the current world average. At the same time, we observe that the preferred $\alpha_{s}$ values vary across the processes. For example, the DIS experiments prefer the lowest value ($\alpha_{s}(M_{Z})=0.1152$). [^5] In the CT25 fit, the majority of DIS experiments prefer $\alpha_{s}(M_{Z})<0.118$, but this preference is especially pronounced for the HERA DIS data at $Q<3.5$ GeV and BCDMS $F_{2}^{p}$ data. Drell–Yan pair and top-quark pair production prefer higher values, 0.1198 and 0.1194, respectively. Inclusive jet production prefers an intermediate value ($\alpha_{s}(M_{Z})=0.1187$). The spread of preferences in Fig. 4 (left) reflects the differing pulls on $\alpha_{s}$ from various experiments and highlights tensions in the data.
When the ksys=2 prescription is applied to the jet data only in Fig. 4 (right), the global minimum of total $\chi^{2}_{R}$ prefers a much lower value,
@@
In this scenario, the ordering of preferred values is markedly different: jet production measurements favor by far the smallest value ($\alpha_{s}(M_{Z})=0.1099$), $t\bar{t}$ production still pulls toward the largest value ($\alpha_{s}(M_{Z})=0.1191$), and DIS ($\alpha_{s}(M_{Z})=0.1146$) and Drell–Yan process ($\alpha_{s}(M_{Z})=0.1187$) lie in-between. Compared to the ksys=1 case, the pull from jet-production measurements to a significantly lower value of $\alpha_{s}$ is accompanied by an equally significant reduction in the jets’ $\chi^{2}$, which reflects strong sensitivity of these measurements to the treatment of correlated systematics.
@@
To further investigate the differences between the ksys=1 and ksys=2 scenarios in the case of inclusive-jet production measurements (in the CT25prel fit), Fig. 5 separately illustrates the $\Delta\chi^{2}$ profiles of individual jet experiments. As before, the curves have $\Delta\chi^{2}=0$ at the best-fit $\alpha_{s}$ of the respective global fit. The ksys=1 case is shown in the Fig. 5 (left), where the DØ Run-2 and ATLAS measurements at 7, 8, and 13 TeV collision energies prefer relatively large $\alpha_{s}$, while the CDF Run-2 and CMS measurements at 7, 8, and 13 TeV prefer smaller values. Their interplay produces a combined $\alpha_{s}$-minimum from jet production at
@@
consistent with the world average.
By contrast, when the ksys=2 treatment is applied to jet data sets in Fig. 5 (right), all of them prefer a much smaller $\alpha_{s}$. The CMS 8 TeV and CDF Run-2 data especially pull toward smaller $\alpha_{s}$ values, compared to the preference of DØ and ATLAS for a somewhat larger $\alpha_{s}$. This produces a much lower central value
@@
that is well below the world average and with a larger uncertainty.
In summary, the comparison between the two different treatments for the correlated systematic uncertainties in inclusive jet production, ksys=1 and ksys=2, demonstrates that these measurements are particularly sensitive to the normalization of systematic errors. [^6] An analogous sensitivity has been noted in the NNPDF precision determination [24] of $\alpha_{s}(M_{Z})$, whereby a fit with a floating $\alpha_{s}$ in the covariance matrix produces a higher value than the final result consistent with the additive prescription (the analog of our ksys=2). As we emphasized, requiring all correlated errors to be normalized according to the ksys=2 (additive) choice may exaggerate the downward pull on $\alpha_{s}$; however, other (theoretical) biases may be present also with the ksys=1 prescription. Given that the major experimental errors in jet production are regarded as multiplicative, this exercise may exaggerate the possible effect, but still serves as a useful reminder of a vulnerability in the standard publication format for systematic errors.
IV Uncertainty estimation: tolerances
IV.1 The $\Delta\chi^{2}=1$ criterion
@@
We first observe that the dependence of the total $\chi^{2}$ observed in our $\alpha_{s}$ scans, such as the one showed by the solid line in Fig. 1, is highly consistent with a parabolic fit, with the logarithmic derivative $d\ln{\Delta\chi^{2}_{R}}/d(\ln{\alpha_{s}-\alpha_{s,0}})$ staying in the range 1.9-2.2 over the interval $\alpha_{s}(M_{Z})=0.110-0.125$ covered in the CT25 scan. Based on the total $\chi^{2}$, the standard deviation is estimated as $\delta_{\textrm{68\%}}\alpha_{s}(M_{Z})\sim 0.0004$ either from the quadratic fit $\chi^{2}=6.695\cdot 10^{6}\ \left(\alpha_{s}(M_{Z})-0.1183\right)^{2}$ or from the condition $\Delta\chi^{2}=1$ applicable for a single-parameter probability distribution [71].
The $\Delta\chi^{2}=1$ uncertainty is interpreted as the frequentist 68% central confidence interval on the $\alpha_{s}$ parameter” determined from imaginary repetitions (replicas) of the same experiments using the same fitting methodology. Each such replica consists of the same selection of data points, which differ only in their random fluctuations within Gaussian uncertainties of the input measurements.
But, this surrogate parameter may differ from the true $\alpha_{s}$ realized in nature if the input data are inconsistent. Such tensions reveal non-negligible data selection uncertainty, which is a part of the model (epistemic) uncertainty.
[…]
When theory perfectly agrees with data, the $\chi^{2}_{E}/N_{\textrm{pt}}$ ratios of individual experiments fluctuate around the expectation of one with a known probability. What is actually seen in all global fits is that the spread of $\chi^{2}_{E}/N_{\textrm{pt}}$ is wider than this normal distribution arising from purely stochastic fluctuations. One can demonstrate this enlargement of discrepancies by considering histograms of effective Gaussian variables $S_{E}$ constructed from $\chi^{2}_{E}$ [39, 16]. While, in the absence of tensions, the distribution of $S_{E}$ is consistent with the standard normal distribution $\mathcal{N}(0,1)$, the empirically obtained histograms of $S_{E}$ for the CT, MSHT, and NNPDF global fits are wider than $\mathcal{N}(0,1)$ [25].
In this and the following section, we compare several strategies to account for the data selection uncertainty. First, we focus on traditionally used methods that increase the input uncertainties or introduce tolerance on the final uncertainty to address incomplete knowledge of the underlying model. Then, we apply Bayesian methods to tackle mutual consistency of the individual $\alpha_{s}$ determinations.
@@
    Section IV.2 applies the Particle Data Group (PDG) prescription for combining inconsistent measurements.
@@
    Sections IV.3 and IV.4 review the global and dynamical tolerance prescriptions; in Sec. IV.5, we explore stability of the dynamic tolerance with respect to clustering of data points.
@@
    Section IV.6 explores the impact of data reclustering on the $\alpha_{s}$ uncertainty.
  4. 4.
    Section V estimates the $\alpha_{s}$ uncertainty using a Bayesian hierarchical model and Bayesian model averaging.
IV.2 The Particle Data Group prescription
[…]
$\chi^{2}(\alpha_{s},\alpha_{s,i},\sigma_{i})=\frac{(\alpha_{s}-\alpha_{s,i})^{2}}{\sigma_{i}^{2}}\ ,$ (11)
with free parameters $\alpha_{s,i}$ and $\sigma_{i}$ denoting the central value and uncertainty. Table 2 lists preferred $\alpha_{s,i}$ and $\sigma_{i}$, as well as their combination given by the weighted average.
[…]
$\alpha_{s,i}(M_{Z})\times 10^{3}$ 115.14 118.62 120.65 118.37
Error $(\sigma_{i}\times 10^{3})$ 0.629 0.743 0.543 0.360
[…]
The combination value $\overline{\alpha}_{s}=0.1184\pm 0.0004$ is consistent with the respective direct determination from total $\chi^{2}$ in Sec. IV.1. Again, the resulting uncertainty is significantly smaller than for the current world average [72]. More importantly, the level of tension between the input data classes is characterized by the large value of $\sum_{i}\chi^{2}(\overline{\alpha}_{s},\alpha_{s,i},\sigma_{i})\approx 44$ for (3-1) degrees of freedom (d.o.f.), indicating a dramatically bad quality of the fit.
As such disagreement is unlikely to occur randomly, the Introduction of the Review of Particle Physics [72] describes a corrective approach that inflates the input uncertainties or, equivalently, defines the standard deviation to correspond to $\Delta\chi^{2}=T^{2}$ with $T^{2}>1$. The error-inflation approach introduces a scale factor $e_{S_{\text{PDG}}}$ that increases the uncertainties on each individual measurement so that the reduced $\chi^{2}/\text{d.o.f.}\approx 1$. This condition is guaranteed by choosing the scale factor to be
$e_{S_{\text{PDG}}}=\sqrt{\chi^{2}/\text{d.o.f.}}\ .$ (12)
$\alpha_{s}=0.1184\pm 0.0017\quad(\text{scaled errors, }e_{S_{\text{PDG}}}\approx 4.7).$ (13)
The PDG prescription rests on the premise that the nominal uncertainties must be increased when tensions suggest possible mismodeling. It is important to note that, as explicitly recommended by the PDG, such a democratic rescaling of all input uncertainties is not advised especially for fits of poor quality, when a more informed procedure should be used. For example, one could try to introduce an additional systematic error contributing only along the parameter direction affected by the tension [Sec. IV.J in Ref. 25].
IV.3 Global and Dynamical Tolerance Criteria
Nevertheless, the PDG prescription captures the broad idea that, when the control of the underlying model is in doubt, a more conservative estimate linked to testing of the models, i.e., the goodness-of-fit criterion in the frequentist paradigm, may be warranted. The tolerance prescriptions developed by the PDF analyses implement this idea.
The simplest global tolerance [78] uniformly associates the 68% CL uncertainty with an increase $\Delta\chi^{2}=T^{2}>1$ along any direction in space of free parameters in the Hessian formalism. This prescription may be too simplistic, as e.g. tensions among experiments are likely to constrain different combinations of parameters in dissimilar ways. Sometimes one experiment may dominate the constraint along an orthogonal direction (an eigenvector), but more often it is a combination of experiments. The dynamic tolerance, which has been developed in several forms, leverages constraints from individual experiments to obtain $T^{2}$ values independently for each orthogonal direction in the parameter space. In this study, we examine tolerance along the direction corresponding to the $\alpha_{s}$ free parameter. We also discuss how the dynamic tolerance relates to the global and $\Delta\chi^{2}=1$ ones under reclustering of input data.
Global tolerance (GT). A common way to introduce either the global or dynamic tolerance is to consider quantiles of the probability density $P(x)$ supported on $\infty<x<\infty$. The quantile $\xi(P,v)$ at cumulative probability $v$ is defined by
@@ IV.3 Global and Dynamical Tolerance Criteria
One might consider $P(x)$ to be the $\chi^{2}$ distribution of the baseline data set. Then, one determines the global tolerance $T^{2}$, or equivalently constructs the 68% CL interval $\alpha_{s,\min}\leq\alpha_{s}(M_{Z})\leq\alpha_{s,\max}$, from the quantile $\xi(\chi^{2}(N_{\textrm{pt}}),0.68)$ of the $\chi^{2}$ distribution with $N_{\textrm{pt}}$ degrees of freedom. This is a more general procedure to compute $T^{2}$ than the PDG-inspired estimate in Sec. IV.2, as it operates directly with $P(x)$ instead of effective data points. In practice, other values of $T$ may be implemented, such as the GT equal to the average of $T$ values along independent parameter directions in the DT prescription.
Dynamical Tolerance (DT). When the quantiles in Eq.(14) are computed for each $\chi^{2}_{E}$ profile of the individual data sets, and then a single uncertainty range is constructed from their combination, we refer to it as the “dynamical tolerance uncertainty” [79, 25]. Suppose the GT applied just to experiment $E$ renders the uncertainty range $\alpha_{s,\min}^{(E)}\leq\alpha_{s}(M_{Z})\leq\alpha_{s,\max}^{(E)}$. The DT interval for the full data set is constructed separately from the upper and lower boundaries for each experiment as
The two-tier CT18 uncertainty. The CT [39, 68] and MSHT [79] groups implement variations of this generic procedure. The global and dynamical tolerance can be combined into a more complex constraint, such as the two-tier tolerance prescription used by CT10, CT14, and CT18 PDFs. In this prescription, the global tolerance with $T^{2}\sim 37$ is combined with the “tier-2” dynamical tolerance at the 68% CL. The goal is of the current study is to bypass this definition, which otherwise would require, in particular, to estimate the global $T^{2}$ through the quantile argument or by another means.
A simplified tolerance. We will approximate both the CT and MSHT procedures by the following simplified prescription based on the effective Gaussian variable $S$.
$S$ represents a type of a $Z$ score, i.e., it is a function of $\chi^{2}$ that is distributed approximately according to $\mathcal{N}(0,1)$ and is easier to work with than the $\chi^{2}$ distribution with $N_{\textrm{pt}}$ degrees of freedom. For each $\chi^{2}$, the corresponding $S(\chi^{2},N_{\textrm{pt}})$ can be computed using either a simpler Fisher’s approximation [80],
@@
IV.4 Properties of the dynamic tolerance
The dynamical tolerance (DT) has proven to be convenient when accounting for constraints on PDF uncertainties from individual experiments. That said, at the detailed level, the DT has a more complicated behavior than the GT. We illustrate these aspects using the results of the $\alpha_{s}$ scans with the CT25 and CT25prel baselines described in Sec. III.
@@ IV.4 Properties of the dynamic tolerance
DT intervals tend to be asymmetric around the central value given by the position of the global minimum for all experiments. This can be explicitly seen from Eqs. (18) and (15), showing that the DT intervals are determined from changes in the probabilities of two most constraining experiments $E$ with respect to the global $\chi^{2}$ minimum. [^8] In the complete the CT or MSHT implementations of the DT formulas, such behavior is achieved by applying the quantile to the ratio $S(\chi^{2}_{E})/S(\chi^{2}_{E,0})$ or $\chi^{2}_{E}/\chi^{2}_{E,0}$, not to $S(\chi^{2}_{E})$ or $\chi^{2}_{E}$ itself. For example, Fig. 6 compares the GT and DT uncertainties in the CT25 NNLO fit for credibility levels indicated on the horizontal axis. Throughout this article, the GT is found from the quantile of the baseline data set according to Eq. (14). The DT is constructed from the quantiles of 49 contributing data sets and has an obvious asymmetry above and below the central $\alpha_{s}(M_{Z})$.
@@
Figure 7 shows the same default GT and DT uncertainties for the 49 data sets as in Fig. 6. In addition, we recomputed the DT 10000 times, every time randomly combining the 49 data sets into 24 clusters in the left subfigure and into 3 clusters in the right one. For every CL value, we thus quantify the uncertainty on the DT uncertainty, communicated in the figure by the filled blue bands containing 6800 instances of the re-clustered DT intervals, as well as the dotted blue bands (envelopes) indicating the most extreme instances among the 10000 re-clustered DT intervals. From Fig. 7, it is clear that the average length and uncertainty of the DT interval depend on reclustering. Fewer clusters result in a DT uncertainty closer to the GT uncertainty, corresponding in this example to a single global data cluster. Conversely, finer clustering reduces the DT uncertainty interval.
@@
Figure 8 summarizes this behavior by showing the DT uncertainty at 68% CL and 90% CL for the CT25prel baseline data comprised of 52 published data sets. In addition to the GT uncertainty, corresponding to one cluster containing all data, we show the DT uncertainty for 52 published data sets, as well as for alternative clusters (both supersets and subsets) detailed in Sec. IV.6.
The figure makes it obvious that the DT uncertainty changes depending on the clustering procedure. Specifically: